If repetition is not allowed then how many numbers between 2000 and 3000 can be formed using the digits from 0 to 7?

Answer by math_tutor2020(3817)


The left most slot (thousands digit) can only be a 2 if we want the number between 2000 and 3000.

The set {0,1,2,3,4,5,6,7} is reduced to {0,1,3,4,5,6,7}
The original set had 8 items in it and the new reduced set has 7 items in it.

For the second slot, there are 7 choices to pick from (pick anything in that reduced set).

Then the third slot has 7-1 = 6 items to pick from if we aren't allowed to repeat digits.

Lastly, the final slot has 6-1 = 5 items to pick from.

Multiply those values: 7*6*5 = 42*5 = 210

Answer: 210